OpenStudy (freckles):

\[\text{ Let } n \text{ be a positive integer with } n \ge 3. \text{ Show } 1989|(n^{n^{n^{n}}}-n^{n^{n}}) .\]

OpenStudy (anonymous):

does \[1989|(n^{n^{n^{n^n}}}-n^{n^{n^n}}) .\] too?

OpenStudy (freckles):

If this helps anyone this comes from http://www.math.muni.cz/~bulik/vyuka/pen-20070711.pdf . A10.

OpenStudy (freckles):

lol I don't know about that I don't know how to do this one honestly

OpenStudy (anonymous):

me neither

OpenStudy (freckles):

@dan815 @rational @kainui @ganeshie8 you guys have any thoughts totally blank right now

OpenStudy (freckles):

disappointed because I thought we were going to get to prove something

OpenStudy (kainui):

Are you sure it's not just such a huge number that it broke wolfram alpha? http://www.wolframalpha.com/input/?i=%285%5E%285%5E%285%5E5%29%29-5%5E%285%5E5%29%29

OpenStudy (anonymous):

i think there are different interpretations of \(n^{n^n}\)

OpenStudy (freckles):

oh i guess it is possible to break wolfram :p

OpenStudy (kainui):

@satellite73 I don't think so, since the alternative is \(n^{n^n}= n^{2n}\) which seems kinda like a waste.

OpenStudy (rational):

n^n^n w/o parenthesis is a tetration i think

OpenStudy (rational):

5^5^5 = 5^(5^5)

OpenStudy (dan815):

@satellite73 here are some useful exponentiation rules you should memorize \[(a^n)^m=a^{nm}\\ a^{n}*a^{m} = a^{n+m}\]

geerky42 (geerky42):

Pretty sure @satellite73 already memorized these rules lol.

OpenStudy (dan815):

how about starting here n^n^n...n-n^n^n...n = 0 mod k k times and k-1 times

OpenStudy (dan815):

ohh hhow about!

OpenStudy (dan815):